OpenStudy (anonymous):

how do i solve 5(x+10)+x

OpenStudy (anonymous):

is it equal to anything?

OpenStudy (mathstudent55):

First, you need to use the distributive property to get rid of parentheses. Then you combine like terms.

OpenStudy (mathstudent55):

Do you know how to use the distributive property?

OpenStudy (anonymous):

You're gifted @FireWolfSpirit I know you can do this.

OpenStudy (nincompoop):

it is not an equation, so what are you trying to solve?

OpenStudy (mathstudent55):

The distributive property of multiplication over addition: \(a(b + c) = ab + ac\) Example: \(5(x + 3) = 5 \cdot x + 5 \cdot 3 = 5x + 15\)

OpenStudy (nincompoop):

5(x+10)+x is only an expression and not an equation

OpenStudy (nincompoop):

if you assumed that the expression equals to zero then you just may be in the right track but without an equation, solving for x is like looking for a miracle.

OpenStudy (anonymous):

@FireWolfSpirit im going to assume your Gaara .... xD

OpenStudy (nincompoop):

5(x+10)+x = 0 THE equal sign is pertinent and must exist for x to be solved.

OpenStudy (nincompoop):

what do you think @wio ?

OpenStudy (anonymous):

You can't solve without an equation. All you can do is simplify.

OpenStudy (nincompoop):

thanks

OpenStudy (anonymous):

hey sorry

OpenStudy (anonymous):

@Ray_08 yes i'm garra @wio thank you and @nincompoop i need to simplify the expression

OpenStudy (anonymous):

I'm thinking:  am i right?